OpenStudy (anonymous):

identify excluded value. y=5/x-4

OpenStudy (anonymous):

set the denominator = 0 and solve for x

OpenStudy (anonymous):

i dont get it

OpenStudy (lgbasallote):

what's your denominator?

OpenStudy (anonymous):

x-4

OpenStudy (lgbasallote):

equate it to zero x - 4 = 0 find x

OpenStudy (anonymous):

???? im so confusewd~!!!!!1

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

\[y=\frac{5}{x-4}\] and you want to know what value you CANNOT substitute for \(x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the only thing you have to worry about is that you are not allowed to divide by 0

OpenStudy (anonymous):

in other words \(\frac{5}{0}\) is undefined, it is not a number

OpenStudy (anonymous):

so you have to only be concerned with what value of \(x\) would make \(x-4\) be zero

OpenStudy (anonymous):

solve this problem by writing \[x-4=0\] and solve for \(x\) in one step if \(x-4=0\) then \(x=4\) so you may NOT have \(x=4\) and that is the excluded value

OpenStudy (anonymous):

ooo ok thanks

OpenStudy (anonymous):

yw